# On face angles of tetrahedra with a given base

### v2

@@ 8.1. On special regions on the cylinder $\operatorname{Cyl}$
$\gamma_{BC}=\mathbb{T}_{BC}^{\,\alpha_{1}}\cap\operatorname{Cyl},\quad\gamma_{AC}=\mathbb{T}_{AC}^{\,\alpha_{2}}\cap\operatorname{Cyl},\quad\gamma_{AB}=\mathbb{T}_{AB}^{\,\alpha_{3}}\cap\operatorname{Cyl}.$ (33)
@@
can be continuously extended to the closed interval $[-\pi/2,\pi/2]$, where $\gamma_{BC}(-\pi/2)=(0,-1,0)$ and $\gamma_{BC}(\pi/2)=(0,1,0)$.
@@
$0\leq z\leq 2\sqrt{\cos(\alpha_{3})\cos(\varphi-\alpha_{3})}$ for $\mbox{for\,\,\,}\varphi\in[\alpha_{3}-\pi/2,\varphi_{0}],$
$0\leq z\leq 2\sqrt{\cos(\alpha_{2})\cos(\varphi+\alpha_{2})}$ for $\mbox{for\,\,\,}\varphi\in[\varphi_{0},-\alpha_{2}+\pi/2],$
@@
where $\varphi\in[-\pi/6,5\pi/6]$. The curves $\gamma_{BC}$ and $\gamma_{AB}$ are obtained by rotating $\gamma_{AC}$ around the third coordinate axis through the angles ±$\pm 2\pi/3$.
@@
[…]
@@ 8.3. On the region inside $\mathbb{BP}$
$\mathbb{R}^{3}\ni(x_{0},y_{0},z_{0}):=(\cos\angle BAC,\cos\angle ABC,\cos\angle ACB\Bigr{)}.$ (34)
@@
We will call it $\Xi(i,j,k)$, where $i=\operatorname{sgn}(x-x_{0})$, $j=(y-y_{0})$, $k=\operatorname{sgn}(z-z_{0})$, and $(x,y,z)$ is any point in a given subset. We are going to mark every $\Xi(i,j,k)$ by the symbol ‘‘$+$’’ (respectively, ‘‘--’’) if $\Xi(i,j,k)$ contains a point in the interior of the set $\operatorname{CFT}$ (respectively, a point in the interior of the set $\mathbb{R}^{3}\setminus\operatorname{CFT}$). If $\Xi(i,j,k)$ contains some points of the interior of $\operatorname{CFT}$ as well as some points of $\mathbb{R}^{3}\setminus\operatorname{CFT}$, then we mark it by symbol ‘‘$\pm$’’. We will do it separately for the cases of acute-angled, right-angled and obtuse-angled base.
By Corollary 3, the point $(x_{0},y_{0},z_{0})$ is situated in the interior of the set $\operatorname{CFT}$ (the closure of $F(\operatorname{GT})$) if the triangle $ABC$ is acute-angled (i.e., $z_{0}$ >0). Therefore, each region $\Xi(i,j,k)$ contains a point in the interior of the set $\operatorname{CFT}$. Moreover, in the acute-angled case we have $F(\mathbb{B}_{-})=\Xi(-,-,-)\bigcap\mathbb{BP}$, therefore, $\Xi(-,-,-)$ should be marked by ‘‘$+$’’. It is easy to see also that
@@
therefore, $\Xi(-,+,+)$, $\Xi(+,-,+)$, $\Xi(+,+,-)$, and $\Xi(+,+,+)$ should also be marked by ‘‘$+$’’. On the other hand, $\Xi(-,-,+)$, $\Xi(-,+,-)$, and $\Xi(+,-,-)$ should be marked by ‘‘$\pm$’’. This observation corresponds to the fact that the images of special regions on the cylinder $\operatorname{Cyl}$ (see Definition 1) are in the subsets $\Xi(-,-,+)$, $\Xi(-,+,-)$, and $\Xi(+,-,-)$, see Subsections 8.1 and 8.2.
@@
$z_{0}>0$ $+$ $\pm$ $\pm$ $+$ $\pm$ $+$ $+$ $+$
$z_{0}=0$ $+$ $\pm$ -- $+$ -- $+$ -- $+$
$z_{0}<0$ $+$ $\pm$ -- $+$ -- $+$ -- $+$
The cases of right-angled and obtuse-angled bases can be considered similarly. The principal feature of these cases is that only $\Omega(-,-,+)$ should be marked by ‘‘$\pm$’’. It is related to the fact that there is only one special region on the cylinder $\operatorname{Cyl}$ for these cases, see Subsection 8.1. All corresponding computations are included in Table 1.
@@ 9. The main construction
$\widetilde{A}$ ::=:= $\Bigl{(}-\cos\angle BAC,\cos\angle ABC,\cos\angle ACB\Bigl{)}\in E_{B}\bigcap E_{C},$
$\widetilde{B}$ ::=:= $\Bigl{(}\cos\angle BAC,-\cos\angle ABC,\cos\angle ACB\Bigl{)}\in E_{A}\bigcap E_{C},$
$\widetilde{C}$ ::=:= $\Bigl{(}\cos\angle BAC,\cos\angle ABC,-\cos\angle ACB\Bigl{)}\in E_{A}\bigcap E_{B}.$
@@
For any given base $ABC$, the closure $\operatorname{CFT}$ of $F(\operatorname{GT})$ is such that $\operatorname{CFT}\subset\mathbb{P}$ and is homeomorphic to a $3$-dimensional closed ball, while the boundary $\operatorname{BFT}$ of $\operatorname{CFT}$ is homeomorphic to a $2$-dimensional sphere $S^{2}$. Moreover, $\operatorname{BFT}\bigcap\mathbb{BP}$ is the union of $F(\mathbb{B}_{-})$, $F(\mathbb{B}_{+})$, $E_{A}$, $E_{B}$, and $E_{C}$, while $\operatorname{BFT}\setminus\mathbb{BP}$ is the union of several connected subsets of the limit solid ellipses $\operatorname{Lim}(A)$, $\operatorname{Lim}(B)$, and $\operatorname{Lim}(C)$, and the images of special regions on the cylinder $\operatorname{Cyl}$.
@@
Since $\operatorname{Cyl}\bigcap\mathbb{B}_{-}=\emptyset$ and $\operatorname{Cyl}\bigcap\mathbb{B}_{+}=\emptyset$, therefore some part of $F(\operatorname{Cyl})$ can be a part of $\operatorname{BFT}$, only if its boundary is the union of some curves in the limit solid ellipses $\operatorname{Lim}(A)$, $\operatorname{Lim}(B)$, and $\operatorname{Lim}(C)$ (see Proposition 10). It should be noted that some part of any of the limit segments $[\widetilde{B},\widetilde{C}]$, $[\widetilde{A},\widetilde{C}]$, $[\widetilde{A},\widetilde{B}]$ may be part of such a boundary (see Proposition 14). Thus, the most important role in this construction is played by curves that are the intersections of the image of the cylinder $\operatorname{Cyl}\bigcap\Pi^{+}$ with the limit solid ellipses $\operatorname{Lim}(A)$, $\operatorname{Lim}(B)$, and $\operatorname{Lim}(C)$.
@@
$\Sigma(\triangle ABC)=F(\Pi^{+})=\operatorname{int}(\operatorname{CFT})\bigcup F(\operatorname{SRT}).$ (35)

### v3

@@ 8.1. On special regions on the cylinder $\operatorname{Cyl}$
(33) $\gamma_{BC}=\mathbb{T}_{BC}^{\,\alpha_{1}}\cap\operatorname{Cyl},\quad\gamma_{AC}=\mathbb{T}_{AC}^{\,\alpha_{2}}\cap\operatorname{Cyl},\quad\gamma_{AB}=\mathbb{T}_{AB}^{\,\alpha_{3}}\cap\operatorname{Cyl}.$
@@
can be continuously extended to the closed interval $\pi/2,\pi/2]$, where $\gamma_{BC}(-\pi/2)=(0,-1,0)$ and $\gamma_{BC}(\pi/2)=(0,1,0)$.
@@
$0\leq z\leq 2\sqrt{\cos(\alpha_{3})\cos(\varphi-\alpha_{3})}$ $\mbox{for\,\,\,}\varphi\in[\alpha_{3}-\pi/2,\varphi_{0}],$
$0\leq z\leq 2\sqrt{\cos(\alpha_{2})\cos(\varphi+\alpha_{2})}$ $\mbox{for\,\,\,}\varphi\in[\varphi_{0},-\alpha_{2}+\pi/2],$
@@
where $\varphi\in[-\pi/6,5\pi/6]$. The curves $\gamma_{BC}$ and $\gamma_{AB}$ are obtained by rotating $\gamma_{AC}$ around the third coordinate axis through the angles $\pm 2\pi/3$.
@@
$\gamma_{AC}(\varphi)=\left(\,\cos\varphi,\,\sin\varphi,\,\sqrt{2\cos\varphi+2\sin\varphi}\,\right),\mbox{\,\,\,where\,\,\,}\varphi\in\left[-\pi/4,3\pi/4\right];$
$\gamma_{BC}(\varphi)=\left(\,\cos\varphi,\,\sin\varphi,\,\sqrt{-\sqrt{3}\cos\varphi+3\sin\varphi}\,\right),\mbox{\,\,\,where\,\,\,}\varphi\in\left[{\pi}/{6},{7\pi}/{6}\right].$
@@ 8.3. On the region inside $\mathbb{BP}$
(34) $\mathbb{R}^{3}\ni(x_{0},y_{0},z_{0}):=(\cos\angle BAC,\cos\angle ABC,\cos\angle ACB\Bigr).$
@@
We will call it $\Xi(i,j,k)$, where $i=\operatorname{sgn}(x-x_{0})$, $j=(y-y_{0})$, $k=\operatorname{sgn}(z-z_{0})$, and $(x,y,z)$ is any point in a given subset. We are going to mark every $\Xi(i,j,k)$ by the symbol ‘‘++’’ (respectively, ‘‘-’’) if $\Xi(i,j,k)$ contains a point in the interior of the set $\operatorname{CFT}$ (respectively, a point in the interior of the set $\mathbb{R}^{3}\setminus\operatorname{CFT}$). If $\Xi(i,j,k)$ contains some points of the interior of $\operatorname{CFT}$ as well as some points of $\mathbb{R}^{3}\setminus\operatorname{CFT}$, then we mark it by symbol $\pm$’’. We will do it separately for the cases of acute-angled, right-angled and obtuse-angled base.
By Corollary 3, the point $(x_{0},y_{0},z_{0})$ is situated in the interior of the set $\operatorname{CFT}$ (the closure of $F(\operatorname{GT})$) if the triangle $ABC$ is acute-angled (i.e., $z_{0}$ >0). Therefore, each region $\Xi(i,j,k)$ contains a point in the interior of the set $\operatorname{CFT}$. Moreover, in the acute-angled case we have $F(\mathbb{B}_{-})=\Xi(-,-,-)\bigcap\mathbb{BP}$, therefore, $\Xi(-,-,-)$ should be marked by ‘‘++’’. It is easy to see also that
@@
therefore, $\Xi(-,+,+)$, $\Xi(+,-,+)$, $\Xi(+,+,-)$, and $\Xi(+,+,+)$ should also be marked by ‘‘++’’. On the other hand, $\Xi(-,-,+)$, $\Xi(-,+,-)$, and $\Xi(+,-,-)$ should be marked by $\pm$’’. This observation corresponds to the fact that the images of special regions on the cylinder $\operatorname{Cyl}$ (see Definition 1) are in the subsets $\Xi(-,-,+)$, $\Xi(-,+,-)$, and $\Xi(+,-,-)$, see Subsections 8.1 and 8.2.
@@
$z_{0}>0$ ++ $\pm$ $\pm$ ++ $\pm$ ++ ++ ++
$z_{0}=0$ ++ $\pm$ - ++ - ++ - ++
$z_{0}<0$ ++ $\pm$ - ++ - ++ - ++
The cases of right-angled and obtuse-angled bases can be considered similarly. The principal feature of these cases is that only $\Omega(-,-,+)$ should be marked by $\pm$’’. It is related to the fact that there is only one special region on the cylinder $\operatorname{Cyl}$ for these cases, see Subsection 8.1. All corresponding computations are included in Table 1.
@@ 9. The main construction
$\widetilde{A}$ $:=$ $\Bigl(-\cos\angle BAC,\cos\angle ABC,\cos\angle ACB\Bigl)\in E_{B}\bigcap E_{C},$
$\widetilde{B}$ $:=$ $\Bigl(\cos\angle BAC,-\cos\angle ABC,\cos\angle ACB\Bigl)\in E_{A}\bigcap E_{C},$
$\widetilde{C}$ $:=$ $\Bigl(\cos\angle BAC,\cos\angle ABC,-\cos\angle ACB\Bigl)\in E_{A}\bigcap E_{B}.$
@@
For any given base $ABC$, the closure $\operatorname{CFP}$ of $F(\operatorname{GT})$ is such that $\operatorname{CFP}\subset\mathbb{P}$ and is homeomorphic to to a $3$-dimensional closed ball, while the boundary $\operatorname{BFP}$ of $\operatorname{CFP}$ is homeomorphic to a $2$-dimensional sphere $S^{2}$. Moreover, $\operatorname{BFP}\bigcap\mathbb{BP}$ is the union of $F(\mathbb{B}_{-})$, $F(\mathbb{B}_{+})$, $\{(1,1,1)\}$, $E_{A}$, $E_{B}$, and $E_{C}$, while $\operatorname{BFP}\setminus\mathbb{BP}$ is the union of several connected subsets of the limit solid ellipses $\operatorname{Lim}(A)$, $\operatorname{Lim}(B)$, and $\operatorname{Lim}(C)$, and the images of special regions on the cylinder $\operatorname{Cyl}$.
@@
Since $\operatorname{Cyl}\bigcap\mathbb{B}_{-}=\emptyset$ and $\operatorname{Cyl}\bigcap\mathbb{B}_{+}=\emptyset$, therefore some part of $F(\operatorname{Cyl})$ can be a part of $\operatorname{BFT}$, only if its boundary is the union of some curves in the limit solid ellipses $\operatorname{Lim}(A)$, $\operatorname{Lim}(B)$, and $\operatorname{Lim}(C)$ (see Proposition 10). It should be noted that some part of any of the limit segments $\widetilde{B},\widetilde{C}$], $\widetilde{A},\widetilde{C}$], $\widetilde{A},\widetilde{B}$] may be part of such a boundary (see Proposition 14). Thus, the most important role in this construction is played by curves that are the intersections of the image of the cylinder $\operatorname{Cyl}\bigcap\Pi^{+}$ with the limit solid ellipses $\operatorname{Lim}(A)$, $\operatorname{Lim}(B)$, and $\operatorname{Lim}(C)$.
@@
(35) $\Sigma(\triangle ABC)=F(\Pi^{+})=\operatorname{int}(\operatorname{CFT})\bigcup F(\operatorname{SRT}).$